# Inference Scaling Laws: An Empirical Analysis of Compute-Optimal Inference for LLM Problem-Solving

### v2

@@ Conclusions.
Our results lead to three main takeaways. First, we find that using a smaller model and generating more tokens in an inference strategy often outperforms using a larger model at a fixed compute budget. This has implications for models deployed in the real world, where inference compute is constrained in various ways. Specifically, it is potentially beneficial to deploy smaller models with more sophisticated inference strategies for better cost-performance tradeoffs. Second, we showed that in the limit of infinite compute (allocated by drawing more samples), sampling-based majority voting strategies inevitably saturate to a distribution that depends on the underlying generation policy. Hence, it is of interest to alter the sampling distribution by designing an alternative inference strategy. Third, we design such an inference strategy–the novel Rebase tree search–and find it is Pareto optimal, in that it achieves the best performance across all tested compute budgets. Notably, it outperforms commonly used weighted majority voting and MCTS methods that have attracted much interest and widespread use. This finding not only shows the strength of Rebase, but also indicates that there is large headroom to improve language model performances via inference-time algorithms.
Limitations.
Our empirical analysis specifically targets mathematical problem-solving. Investigating the inference scaling laws and compute-optimal inference strategies for tasks beyond mathematical problem-solving would be a valuable direction for future research. Additionally, we mainly evaluate the proposed Rebase on the GSM8K and MATH500 datasets. We speculate that the Rebase algorithm, which assumes access only to a function that assigns scores to nodes, will be effective in tasks beyond those studied here.
@@ Appendix C Hyper-parameters
Finetuning
All the hyperparameters for model fine-tuning can be found in Table 2. We preprocess the MetaMath Dataset to make the solutions in a stepwise format.
Inference
@@
For all the inference strategies, the temperature of the LLM is set to $1.0$. Max tokens for the output is $1024$ and max tokens for one step is $256$. For REBASE, we chose the balance temperature (the softmax temperature in the REBASE algorithm) as $T_{b}=0.1$. For MCTS, we set $C$ in the UCT value as 1 and we expand $4,8,16$ children for the root, 2 children for other selected nodes with total 32, 64, 128 expansions respectively.
Appendix D Supplementary Figures and Tables
@@
Figure 8: The inference computation scaling laws of different models for the problem-solving error rate on MATH test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, M.V. refer to majority voting.
@@
Figure 9: The inference computation scaling laws of different models for the problem-solving error rate on GSM8K test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, M.V. refer to majority voting.
@@
In this section, we additionally include experimental results on the majority voting method, along with its comparison with weighted majority voting (Fig. 8, 9 ,10, 11). The experiments show that although the gap between majority voting and weighted majority voting on sampling is huge. This gap becomes much smaller if we apply Rebase. This phenomenon can be caused by the selection ability of tree search like Rebase. Once Rebase already samples solutions with high rewards, conducing weighted majority voting gains less since the sampled solutions may all have relatively high and stable rewards compared with those of sampling.
@@
Figure 10: The inference computation scaling laws of different models for the problem-solving error rate on MATH test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, M.V. and W.M. refer to majority voting and weighted majority, respectively.
@@
Figure 11: The inference computation scaling laws of different models for the problem-solving error rate on GSM8K test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, M.V. and W.M. refer to majority voting and weighted majority, respectively.

### v3

@@ 5 Conclusions
Our results lead to three main takeaways. First, we find that using a smaller model and generating more tokens in an inference strategy often outperforms using a larger model at a fixed compute budget. This has implications for models deployed in the real world, where inference compute is constrained in various ways. Specifically, it is potentially beneficial to deploy smaller models with more sophisticated inference strategies for better cost-performance trade-off. Second, we show that in the limit of infinite compute (allocated by drawing more samples), sampling-based majority voting strategies inevitably saturate to a distribution that depends on the underlying generation policy. Hence, it is of interest to alter the sampling distribution by designing an alternative inference strategy. Third, we design such an inference strategy–the novel Rebase tree search–and find it is Pareto optimal, in that it achieves the best performance across all tested compute budgets. Notably, it outperforms commonly used weighted majority voting and MCTS methods that have attracted much interest and widespread use. This finding not only shows the strength of Rebase, but also indicates that there is large headroom to improve language model performances via inference-time algorithms.
Acknowledgment
Zhiqing Sun acknowledges the support of the Google Fellowship. Sean Welleck thanks NSF SCALE (NSF DMS 2134012) and Convergent Research.
@@ Appendix C Hyper-parameters
Finetuning.
All the hyperparameters for model fine-tuning can be found in Tab. 2. We preprocess the MetaMath Dataset to make the solutions in a stepwise format.
@@
Inference.
@@
For all the inference strategies, the temperature for LLM token generation is set to $1.0$. Max tokens for the output is $1024$ and max tokens for one step is $256$. For Rebase, we set the balance temperature (i.e., the parameter $T_{b}$ in Eq. (1)) to $0.1$. For MCTS, we set $C$ in the UCT value to 1 and we expand $4,8,16$ children for the root, 2 children for other selected nodes with total 32, 64, 128 expansions respectively. New expanded nodes will be assigned values by the PRM, and then backpropagate the $Q$ values through the process described in the last section.
Appendix D Additional experimental results
@@
Figure 9: The inference scaling laws of different models for the problem-solving error rate on MATH test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, M.V. refer to majority voting.
@@
Figure 10: The inference scaling laws of different models for the problem-solving error rate on GSM8K test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, M.V. refer to majority voting.
D.1 Majority voting experiment results
@@
In this section, we additionally include experimental results on the majority voting method, along with its comparison with weighted majority voting (Fig. 9, 10 ,11, 12). The experiments show that although the gap between majority voting and weighted majority voting on sampling is huge. This gap becomes much smaller if we apply Rebase. This phenomenon can be caused by the selection ability of tree search like Rebase. Once Rebase already samples solutions with high rewards, conducing weighted majority voting gains less since the sampled solutions may all have relatively high and stable rewards compared with those of sampling.
@@
Figure 11: The inference scaling laws of different models for the problem-solving error rate on MATH test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, M.V. and W.M. refer to majority voting and weighted majority, respectively.
@@
Figure 12: The inference scaling laws of different models for the problem-solving error rate on GSM8K test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, M.V. and W.M. refer to majority voting and weighted majority, respectively.
D.2 Additional experiments on Llama3 models
We conduct additional experiments with Llama3-8B-Instruct (Dubey et al., 2024) model on MATH and GSM8K datasets, as shown in Fig. 13. Results for code generation task MBPP are presented (Austin et al., 2021) in Tab. 3. These experiments demonstrate that our conclusions generalize to the Llama3 architecture and coding tasks, confirming that increased computational effort improves performance until saturation is reached, and Rebase reaches the optimal performance-compute trade-off.
In mathematical reasoning tasks, Rebase consistently outperforms the sampling approach across different answer selection strategies, including best-of-$n$, majority voting, and weighted voting. The highest performance on each dataset is achieved using Rebase. Specifically, on GSM8K, Rebase combined with weighted majority voting using 128 samples achieves an accuracy of $90.2\%$, surpassing the best accuracy of $89.7\%$ obtained by the sampling method with 256 samples using the best-of-$n$ strategy. Similarly, on MATH, Rebase with weighted majority voting using 128 samples achieves an accuracy of $47.4\%$, significantly outperforming the sampling method’s best accuracy of $41.9\%$ with 256 samples using best-of-$n$.
For the code generation task MBPP, we analyze scaling behavior and compute-optimal inference through pass rate evaluation. The results confirm that Rebase is more compute-efficient than sampling. This advantage can be attributed to the use of a reward model that evaluates partial code solutions. By conducting one iteration of REBASE, our method prunes suboptimal partial solutions while encouraging exploration of promising ones, thereby enhancing computational efficiency and solution quality.
Figure 13: GSM8K (left) and MATH (right) inference scaling across inference strategies and models (lower is better). The tested model is Llama3-instruct-8B. In the legend, M.V., W.M., and BoN refer to majority voting, weighted majority, and best-of-$n$, respectively.
Table 3: Zero shot pass rates of Sampling and Rebase on MBPP code generation task.
# Samples Sampling FLOPs Sampling Pass Rate Rebase FLOPs Rebase Pass Rate
8 $8\times 10^{12}$ 63 $8.3\times 10^{12}$ 69.6
16 $16\times 10^{12}$ 69.4 $17.5\times 10^{12}$ 72.4
32 $32\times 10^{12}$ 72.4 $34.9\times 10^{12}$ 75.8
64 $64\times 10^{12}$ 79 $69.2\times 10^{12}$ 81.4
D.3 Comparison of different strategies across different models
We show the accuracy of different strategies under a specific compute budget in Tab. 4.
Table 4: Accuracy of different inference configurations under a specific compute budget. MV, BoN and WV denote Majority Voting, best-of-$n$ and Weighted Voting, respectively.
# Samples MATH FLOPs GSM8K FLOPs MATH500 GSM8K
Mistral-7B
Greedy 1 $3.4\times 10^{12}$ $2.3\times 10^{12}$ 28.6 77.9
Sampling + MV 32 $109.2\times 10^{12}$ $72.6\times 10^{12}$ 36.1 85.7
Sampling + BoN 32 $109.2\times 10^{12}$ $72.6\times 10^{12}$ 40.3 89.4
Sampling + WV 32 $109.2\times 10^{12}$ $72.6\times 10^{12}$ 39.7 89.1
REBASE + MV 32 $136.2\times 10^{12}$ $78.9\times 10^{12}$ 44.1 88.8
REBASE + BoN 32 $136.2\times 10^{12}$ $78.9\times 10^{12}$ 45.4 89.4
REBASE + WV 32 $136.2\times 10^{12}$ $78.9\times 10^{12}$ 45.0 89.8
Llemma-7B
Greedy 1 $3.92\times 10^{12}$ $2.3\times 10^{12}$ 30.0 68.5
Sampling + MV 32 $125.4\times 10^{12}$ $73.9\times 10^{12}$ 41.0 80.0
Sampling + BoN 32 $125.4\times 10^{12}$ $73.9\times 10^{12}$ 41.7 85.6
Sampling + WV 32 $125.4\times 10^{12}$ $73.9\times 10^{12}$ 43.5 85.4
REBASE + MV 32 $148.0\times 10^{12}$ $82.6\times 10^{12}$ 46.1 86.1
REBASE + BoN 32 $148.0\times 10^{12}$ $82.6\times 10^{12}$ 44.1 86.9
REBASE + WV 32 $148.0\times 10^{12}$ $82.6\times 10^{12}$ 46.8 87.3
Llama3-Instruct-8B
Greedy 1 $3.84\times 10^{12}$ $2.28\times 10^{12}$ 29.6 79.0
Sampling + MV 32 $122.9\times 10^{12}$ $73.2\times 10^{12}$ 35.4 84.6
Sampling + BoN 32 $122.9\times 10^{12}$ $73.2\times 10^{12}$ 39.7 88.5
Sampling + WV 32 $122.9\times 10^{12}$ $73.2\times 10^{12}$ 39.5 88.6
REBASE + MV 32 $172.8\times 10^{12}$ $79.3\times 10^{12}$ 45.2 88.3
REBASE + BoN 32 $172.8\times 10^{12}$ $79.3\times 10^{12}$ 45.5 88.7
REBASE + WV 32 $172.8\times 10^{12}$ $79.3\times 10^{12}$ 43.7 89.1
Llemma-34B
Greedy 1 $19.0\times 10^{12}$ $11.2\times 10^{12}$ 33.0 78.4
Sampling + MV 8 $152.3\times 10^{12}$ $89.7\times 10^{12}$ 39.9 84.3
Sampling + BoN 8 $152.3\times 10^{12}$ $89.7\times 10^{12}$ 40.4 86.7
Sampling + WV 8 $152.3\times 10^{12}$ $89.7\times 10^{12}$ 41.0 86.0
REBASE + MV 8 $176.8\times 10^{12}$ $98.7\times 10^{12}$ 43.9 86.1
REBASE + BoN 8 $176.8\times 10^{12}$ $98.7\times 10^{12}$ 43.6 86.9
REBASE + WV 8 $176.8\times 10^{12}$ $98.7\times 10^{12}$ 42.9 86.9